How would I figure out: Find an equation for a line with the given properties:
Equation for a line parallel to y= -2x containing the point (7,8)

thank you!

Answer by rfer(16322)

8=-2*7+b
8=-14+b
8+14=b
b=22
-----------
y=-2x+22